### v1

# Combinatorial interpretation of a truncated identity of Gauss

Abstract. In 2012, Andrews and Merca proved a truncated theorem on Euler’s pentagonal number theorem, which opened up a new study on truncated theta series. In particular, some results on a truncated identity of Guass have been proved. In this article, we provide a new combinatorial interpretation of the truncated identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
Keywords: truncated identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
AMS Classifications: 05A17, 11P84
1 Introduction
A partition $\pi$ of a positive integer $n$ is a finite non-increasing sequence of positive integers $\pi=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ such that $\pi_{1}+\pi_{2}+\cdots+\pi_{m}=n$. The empty sequence forms the only partition of zero. The $\pi_{i}$ are called the parts of $\pi$. Let $p(n)$ denote the number of partitions of $n$. The generating function of $p(n)$ was given by Euler
$\sum_{n=0}^{\infty}p(n)q^{n}=\frac{1}{(q;q)_{\infty}}.$
Here and in the sequel, we assume that $|q|<1$ and employ the standard notation:
$(a;q)_{\infty}=\prod_{i\geq 0}(1-aq^{i}),$
$(a;q)_{n}=\frac{(a;q)_{\infty}}{(aq^{n};q)_{\infty}},$
and
${M\brack N}=\left\{\begin{array}[]{ll}\frac{(q;q)_{M}}{(q;q)_{N}(q;q)_{M-N}},&\text{if }M\geq N\geq 0,\\ 0,&\text{otherwise.}\end{array}\right$.
One of the well-known theorems in the partition theory is Euler’s pentagonal number theorem:
$(q;q)_{\infty}=\sum_{n=0}^{\infty}(-1)^{n}q^{n(3n+1)/2}(1-q^{2n+1}).$ (1.1)
In 2012, Andrews and Merca [2] considered a truncated version of (1.1) and obtained that for $k\geq 1$,
$\frac{1}{(q;q)_{\infty}}\sum_{j=0}^{k-1}(-1)^{j}q^{j(3j+1)/2}(1-q^{2j+1})=1+(-1)^{k-1}\sum_{n=1}^{\infty}\frac{q^{k(k-1)/2+(k+1)n}}{(q;q)_{n}}{{n-1}\brack{k-1}}$,
from which they deduced the following theorem for the partition function $p(n)$.
Theorem $1.$
[2, Theorem 1.1] For $n,k\geq 1$,
$(-1)^{k-1}\sum_{j=0}^{k-1}(-1)^{j}\left(p(n-j(3j+1)/2)-p(n-j(3j+5)/2-1)\right)=M_{k}(n),$
where $M_{k}(n)$ is the number of partitions of $n$ in which $k$ is the least integer that is not a part and there are more parts >k>k than there are <k<k.
Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Guass [1, p. 23]:
$1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}=\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}.$ (1.2)
Inspired by Andrews and Merca’s work, Guo and Zeng [9] established the truncated version of (1.2) in 2013. They proved that for $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k}\sum_{n=k+1}^{\infty}\frac{(-q;q)_{k}(-1;q)_{n-k}q^{(k+1)n}}{(q;q)_{n}}{{n-1}\brack{k}}$. (1.3)
An overpartition, introduced by Corteel and Lovejoy [7], is a partition such that the last occurrence of a number may be overlined. For example, there are fourteen overpartitions of 44.
$(4),(\overline{4}),(3,1),(\overline{3},1),(3,\overline{1}),(\overline{3},\overline{1}),(2,2),(2,\overline{2}),$
$(2,1,1),(\overline{2},1,1),(2,1,\overline{1}),(\overline{2},1,\overline{1}),(1,1,1,1),(1,1,1,\overline{1}).$
We impose the following order on the parts of an overpartition.
$\overline{1}<1<\overline{2}<2<\cdots.$ (1.4)
Let $\overline{p}(n)$ be the number of overpartitions of $n$. The generating function of $\overline{p}(n)$ is
$\sum_{n=0}^{\infty}\overline{p}(n)q^{n}=\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}.$
The following corollary is an immediate consequence of (1.3).
Corollary 1.2.
[9, Corollary 1.2] For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)\geq 0,$ (1.5)
with strict inequality if $n\geq(k+1)^{2}$. For example,
$\overline{p}(n)-2\overline{p}(n-1)\leq 0,\text{ with strict inequality if }n\geq 4.$ (1.6)
A combinatorial proof of (1.6) was given by Guo and Zeng [9]. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
In 2018, Andrews and Merca [6] provided the following version of (1.3). For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(1-q^{k+j+1})(q^{k+j+2};q)_{\infty}},$
from which they deduced the following combinatorial interpretation of the sum in (1.5). The combinatorial proof of the following theorem was given by Ballantine, Merca, Passary and Yee [5].
Theorem 1.3.
[6, Corollary 8] For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=\overline{M}_{k}(n),$
where $\overline{M}_{k}(n)$ is the number of overpartitions of $n$ in which the first part larger than $k$ appears at least $k+1$ times.
In this article, we will present a new combinatorial interpretation of the sum in (1.5) by using the minimal excludant non-overlined part of an overpartition. The minimal excludant of a partition was introduced by Grabner and Knopfmacher [8] under the name “smallest gap”. Recently, Andrews and Newman [3] undertook a combinatorial study of the minimal excludant of a partition. The minimal excludant of a partition $\pi$ is the smallest positive integer that is not a part of $\pi$, denoted $mex(\pi)$. There is a generalization of the minimal excludant of a partition. In [4], Andrews and Newman defined $mex_{A,a}(\pi)$ to be the smallest positive integer congruent to $a$ modulo AA that is not a part of $\pi$. In [10], Yang and Zhou extended the definition of $mex_{A,a}(\pi)$ to overpartitions.
Definition 1.4.
[10, Definition 4.1] Let $\overline{mes}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo AA that is not a non-overlined part in the overpartition $\pi$. For $n\geq 1$, define $op_{A,a}(n)$ (resp. $\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{A,a}(\pi)$ is congruent to $a$ (resp. $a+A$) modulo $2A$.
By definition, it is easy to get that for $n\geq 1$,
$op_{A,a}(n)+\overline{op}_{A,a}(n)=\overline{p}(n).$ (1.7)
For example, we have
$\pi$ $(4)$ $(\overline{4})$ $(3,1)$ $(\overline{3},1)$ $(3,\overline{1})$ $(\overline{3},\overline{1})$ $(2,2)$ $(2,\overline{2})$
$\overline{mes}_{2,1}(\pi)$ 11 11 55 33 11 11 11 11
$\pi$ $(2,1,1)$ $(\overline{2},1,1)$ $(2,1,\overline{1})$ $(\overline{2},1,\overline{1})$ $(1,1,1,1)$ $(1,1,1,\overline{1})$
$\overline{mes}_{2,1}(\pi)$ 33 33 33 33 33 33
Hence, $op_{2,1}(4)=7$ and $\overline{op}_{2,1}(4)=7$.
For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{2,1}(\pi)\geq 2k+1$ and $\overline{mes}_{2,1}(\pi)\equiv 2k+1\pmod{4}$. By definition, we have
$op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$ (1.8)
We give the following combinatorial interpretation of the sum in (1.5).
Theorem 1.5.
For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=2op_{2,1}(n,k+1).$
The following truncated version of (1.2) immediately follows from Theorem 1.5.
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).$
This article is organized as follows. In Section 2, we give a combinatorial proof of (1.6). Then, we show Theorem 1.5 in Section 3.
2 Combinatorial proof of (1.6)
Clearly, in order to prove (1.6), it is equivalent to showing that for $n\geq 1$,
$\overline{p}(n)/2\leq\overline{p}(n-1),\text{ with strict inequality if }n\geq 4.$ (2.1)
To do this, we introduce three sets of overpartitions. Bear in mind that the parts in an overpartition are ordered as in (1.4).
For $n\geq 1$, let $\mathcal{A}(n)$ be the set of overpartitions of $n$ such that the smallest part is non-overlined. For an overpartition $\pi$ in $\mathcal{A}(n)$, if we change the smallest part to an overlined part, then we get an overpartition of $n$ such that the smallest part is overlined, and vice versa. This implies that the number of overpartitions in $\mathcal{A}(n)$ is $\overline{p}(n)/2$.
For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than 11 in $\lambda$ is greater than or equal to two plus the number of parts 11 in $\lambda$.
For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than 11 in $\lambda$ is less than two plus the number of parts 11 in $\lambda$.
To give a combinatorial proof of (2.1), it suffices to show the following theorem.
Theorem 2.1.
There exists a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$ and $\mathcal{C}(n-1)$ is nonempty for $n\geq 4$.
Proof. For $n\geq 1$, let $\pi=(\pi_{1},\ldots,\pi_{m-1},\pi_{m})$ be an overpartition in $\mathcal{A}(n)$. By definition, we know that $\pi_{m}$ is a non-overlined part, which implies that $\overline{1}$ does not appear in $\pi$. Assume that $\pi_{m}=t$, we consider the following two cases.
Case 1: $t=1$. In this case, we set $\lambda=(\pi_{1},\ldots,\pi_{m-1})$.
Case 2: $t\geq 2$. In this case, we set
$\lambda=(\pi_{1},\ldots,\pi_{m-1},\underbrace{1,\ldots,1}_{(t-2)^{\prime}\text{s}},\overline{1}).$
In either case, we get an overpartition $\lambda$ in $\mathcal{B}(n)$. Obviously, the process above is reversible. Now, we have built a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$.
It remains to show that $\mathcal{C}(n-1)$ is nonempty for $n\geq 4$. It can be checked that for $n\geq 4$,
$(\overline{2},\underbrace{1,\ldots,1}_{(n-4)^{\prime}\text{s}},\overline{1})$
is an overpartition in $\mathcal{C}(n-1)$. This completes the proof.
3 Proof of Theorem 1.5
The objective of this section is to give a proof of Theorem 1.5. To do this, we need the following lemma.
Lemma 3.1.
For $n\geq 1$ and $j\geq 1$,
$\overline{p}(n-j^{2})=op_{2,1}(n,j)+op_{2,1}(n,j+1),$
Proof. For an overpartition $\pi$ of $n-j^{2}$, if we add the parts $1,3,\ldots,2j-1$ to $\pi$, then we get an overpartition $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$, and vice versa. It implies that $\overline{p}(n-j^{2})$ equals the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$.
On the other hand, it follows from definition that $op_{2,1}(n,j)+op_{2,1}(n,j+1)$ is also the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$. This completes the proof.
In [10], Yang and Zhou obtained the following theorem.
Theorem 3.2.
[10, Theorem 4.2] For $n\geq 1$, we have
$op_{2,1}(n)=\overline{p}(n)/2.$
Combining with (1.7) and (1.8), we can get the following lemma. We also give a proof of the following lemma with the aid of Lemma 3.1.
Lemma $3.$
For $n\geq 1$,
$\overline{p}(n)=2op_{2,1}(n,1).$
Proof. Divide both sides of (1.2) by $\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}$, we can get
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}\right)=1.$ (3.1)
Comparing the coefficients of $q^{n}$ on the both sides of (3.1), we have
$\overline{p}(n)+2\sum_{j=1}^{\infty}(-1)^{j}\overline{p}(n-j^{2})=0.$
Combining with Lemma 3.1, we get
$\overline{p}(n)=-2\sum_{j=1}^{\infty}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)=2op_{2,1}(n,1).$
This completes the proof.
Now, we are in a position to give a proof of Theorem 1.5.
Proof of Theorem 1.5. By Lemmas 3.1 and 3.3, we have
$\quad(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)$
$=(-1)^{k}\left(2op_{2,1}(n,1)+2\sum_{j=1}^{k}(-1)^{j}(op_{2,1}(n,j)+op_{2,1}(n,j+1))\right)$
$=(-1)^{k}(-1)^{k}2op_{2,1}(n,k+1)$
$=2op_{2,1}(n,k+1),$
and thus the proof is complete.