### v3

# Combinatorial interpretations of truncated versions of a identity of Gauss

Abstract. In 2012, Andrews and Merca proved a truncated theorem on Euler’s pentagonal number theorem, which opened up a new study on truncated theta series. In particular, some truncated versions of a identity of Gauss have been proved. In this article, we provide new combinatorial interpretations of the truncated versions of the identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
Keywords: truncated versions, combinatorial interpretations, a identity of Gauss, the minimal excludant integer, overpartitions
AMS Classifications: 05A17, 11P84
1 Introduction
A partition $\pi$ of a positive integer $n$ is a finite non-increasing sequence of positive integers $\pi=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ such that $\pi_{1}+\pi_{2}+\cdots+\pi_{m}=n$. The empty sequence forms the only partition of zero. The $\pi_{i}$ are called the parts of $\pi$. Let $p(n)$ denote the number of partitions of $n$. The generating function of $p(n)$ was given by Euler.
$\sum_{n=0}^{\infty}p(n)q^{n}=\frac{1}{(q;q)_{\infty}}.$
Here and in the sequel, we assume that $|q|<1$ and employ the standard notation:
$(a;q)_{\infty}=\prod_{i=0}^{\infty}(1-aq^{i}),$
$(a;q)_{n}=\frac{(a;q)_{\infty}}{(aq^{n};q)_{\infty}},$
and
${M\brack N}=\left\{\begin{array}[]{ll}\frac{(q;q)_{M}}{(q;q)_{N}(q;q)_{M-N}},&\text{if }M\geq N\geq 0,\\ 0,&\text{otherwise.}\end{array}\right$.
One of the well-known theorems in the partition theory is Euler’s pentagonal number theorem:
$(q;q)_{\infty}=\sum_{n=0}^{\infty}(-1)^{n}q^{n(3n+1)/2}(1-q^{2n+1}).$ (1.1)
In 2012, Andrews and Merca [2] considered a truncated version of (1.1) and obtained that for $k\geq 1$,
$\frac{1}{(q;q)_{\infty}}\sum_{j=0}^{k-1}(-1)^{j}q^{j(3j+1)/2}(1-q^{2j+1})=1+(-1)^{k-1}\sum_{n=1}^{\infty}\frac{q^{k(k-1)/2+(k+1)n}}{(q;q)_{n}}{{n-1}\brack{k-1}}$,
from which they deduced the following theorem for the partition function $p(n)$.
Theorem 1.1.
[2, Theorem 1.1] For $n,k\geq 1$,
$(-1)^{k-1}\sum_{j=0}^{k-1}(-1)^{j}\left(p(n-j(3j+1)/2)-p(n-j(3j+5)/2-1)\right)=M_{k}(n),$
where $M_{k}(n)$ is the number of partitions of $n$ in which $k$ is the least integer that is not a part and there are more parts >k>k than there are <k<k.
Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Gauss [1, (2.2.12)]:
$1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}=\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}.$ (1.2)
Inspired by Andrews and Merca’s work, Guo and Zeng [10] established the truncated version of (1.2) in 2013. They proved that for $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k}\sum_{n=k+1}^{\infty}\frac{(-q;q)_{k}(-1;q)_{n-k}q^{(k+1)n}}{(q;q)_{n}}{{n-1}\brack{k}}$. (1.3)
An overpartition, introduced by Corteel and Lovejoy [8], is a partition such that the last occurrence of a number may be overlined. For example, there are fourteen overpartitions of $4$.
$(4),(\overline{4}),(3,1),(\overline{3},1),(3,\overline{1}),(\overline{3},\overline{1}),(2,2),(2,\overline{2}),$
$(2,1,1),(\overline{2},1,1),(2,1,\overline{1}),(\overline{2},1,\overline{1}),(1,1,1,1),(1,1,1,\overline{1}).$
We impose the following order on the parts of an overpartition.
$\overline{1}<1<\overline{2}<2<\cdots.$ (1.4)
Let $\overline{p}(n)$ be the number of overpartitions of $n$. The generating function of $\overline{p}(n)$ is
$\sum_{n=0}^{\infty}\overline{p}(n)q^{n}=\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}.$
The following corollary is an immediate consequence of (1.3).
Corollary 1.2.
[10, Corollary 1.2] For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)\geq 0,$ (1.5)
with strict inequality if $n\geq(k+1)^{2}$. For example,
$\overline{p}(n)-2\overline{p}(n-1)\leq 0,\text{ with strict inequality if }n\geq 4.$ (1.6)
A combinatorial proof of (1.6) was given by Guo and Zeng [10]. In 2018, Andrews and Merca [3] provided the following truncated version of (1.2). For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$
from which they deduced the following combinatorial interpretation of the sum in (1.5). The combinatorial proof of the following theorem was given by Ballantine, Merca, Passary and Yee [7].
Theorem 1.3.
[3, Corollary 8] For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=\overline{M}_{k}(n),$
where $\overline{M}_{k}(n)$ is the number of overpartitions of $n$ in which the first part larger than $k$ appears at least $k+1$ times.
The minimal excludant of a partition was introduced by Grabner and Knopfmacher [9] under the name “smallest gap”. Recently, Andrews and Newman [4] undertook a combinatorial study of the minimal excludant of a partition. The minimal excludant of a partition $\pi$ is the smallest positive integer that is not a part of $\pi$, denoted $mex(\pi)$. In [6], Ballantine and Merca proved some inequalities involving the minimal excludant of a partition. They also posed the following conjecture.
Conjecture 1.4.
[6, Conjecture 1] For $n,k\geq 1$,
$\sum_{j=-\infty}^{\infty}(-1)^{j}\overline{M}_{k}(n-j(3j-1))\geq 0,$
with strict inequality if $n\geq(k+1)^{2}$.
Conjecture 1.4 was settled by Yao in [18]. Furthermore, Yao showed that for $\ell\geq 1$,
$\sum_{n=0}^{\infty}\sum_{j=-\infty}^{\infty}(-1)^{j}\overline{M}_{k}(n-\ell j(3j-1)/2)q^{n}=2\frac{(q^{\ell};q^{\ell})_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}\frac{q^{(k+2j+1)^{2}}(1-q^{2k+4j+3})}{(q;q^{2})_{\infty}}.$ (1.7)
In [10], Guo and Zeng conjectured a stronger inequality than (1.5).
Conjecture 1.5.
[10, (6.4)] For $n,k\geq 1$,
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})\geq 0,$ (1.8)
with strict inequality if $n\geq k^{2}$.
Conjecture 1.5 was proved independently by Mao [12] and Yee [19] and then reconfirmed later by Wang and Yee [13]. In [16], Xia, Yee and Zhao proved the following inequality which implies (1.8).
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k(k+1))\geq 0.$
In [11], Li obtained the following truncated version of (1.2). For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=-k}^{k-1}(-1)^{j}q^{j^{2}}=1+(-1)^{k-1}\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$ (1.9)
based on which Li [11] presented a combinatorial interpretation of the sum in (1.8).
Theorem 1.6.
[11, Theorem 1.2] For $n,k\geq 1$,
$(-1)^{k-1}\sum_{j=-k}^{k-1}(-1)^{j}\overline{p}(n-j^{2})=\overline{N}_{k}(n),$
where $\overline{N}_{k}(n)$ is the number of overpartitions of $n$ in which a part of size $k$ has to be overlined, the smallest part >k1>k-1 appears exactly $k$ times and it cannot be overlined.
In [14, 15], Wang and Xiao also obtained (1.9) and gave a combinatorial interpretation of the sum in (1.8). Li [11] generalized the sums in (1.5) and (1.8).
Theorem 1.7.
[11, Corollary 5.3] For integers $m$ and $k$ with $m\leq k$,
$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})\geq 0.$ (1.10)
2 Main results
In this section, we will list the results obtained in this article. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
There is a generalization of the minimal excludant of a partition. In [5], Andrews and Newman defined $mex_{A,a}(\pi)$ to be the smallest positive integer congruent to $a$ modulo $A$ that is not a part of $\pi$. In [17], Yang and Zhou extended the definition of $mex_{A,a}(\pi)$ to overpartitions.
Definition 2.1.
[17, Definition 4.1] Let $\overline{mex}_{A,a}(\pi)$ be the smallest positive integer congruent to $a$ modulo $A$ that is not a non-overlined part in the overpartition $\pi$. For $n\geq 1$, define $op_{A,a}(n)$ (resp. $\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{A,a}(\pi)$ is congruent to $a$ (resp. $a+A$) modulo $2A$.
By definition, it is easy to get that for $n\geq 1$,
$op_{A,a}(n)+\overline{op}_{A,a}(n)=\overline{p}(n).$ (2.1)
For example, we have
$\pi$ $(4)$ $(\overline{4})$ $(3,1)$ $(\overline{3},1)$ $(3,\overline{1})$ $(\overline{3},\overline{1})$ $(2,2)$ $(2,\overline{2})$
$\overline{mex}_{2,1}(\pi)$ 11 11 55 33 11 11 11 11
$\pi$ $(2,1,1)$ $(\overline{2},1,1)$ $(2,1,\overline{1})$ $(\overline{2},1,\overline{1})$ $(1,1,1,1)$ $(1,1,1,\overline{1})$
$\overline{mex}_{2,1}(\pi)$ 33 33 33 33 33 33
Hence, $op_{2,1}(4)=7$ and $\overline{op}_{2,1}(4)=7$.
Yang and Zhou [17] obtained the following theorem.
Theorem 2.2.
[17, Theorem 4.2] For $n\geq 1$, we have
$op_{2,1}(n)=\overline{p}(n)/{2}.$ (2.2)
For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{2,1}(\pi)\geq 2k+1$ and $\overline{mex}_{2,1}(\pi)\equiv 2k+1\pmod{4}$. By definition, we have
$op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$
Combining with (2.1) and (2.2), we can get the following theorem.
Theorem 2.3.
For $n\geq 1$,
$op_{2,1}(n,1)=\overline{p}(n)/{2}.$
In this article, we will give another proof of Theorem 2.3. Then, we will present a combinatorial interpretation of the sum in (1.10) from the point of view of the minimal excludant non-overlined part of an overpartition.
Theorem 2.4.
For integers $m$ and $k$ with $m\leq k$,
$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,a)+(-1)^{m+k}op_{2,1}(n,b+1)\text{ if }mk>0,$ (2.3)
and
$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,a+1)+(-1)^{m+k}op_{2,1}(n,b+1)\text{ if }mk\leq 0,$ (2.4)
where $a=\text{min}\{|m|,|k|\}$ and $b=\text{max}\{|m|,|k|\}$.
Setting m=k<0m=-k<0 and $m=-k+1\leq 0$ in (2.4), we can get new combinatorial interpretations of the sums in (1.5) and (1.8) respectively.
Corollary 2.5.
For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=2op_{2,1}(n,k+1),$ (2.5)
and
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})=op_{2,1}(n,k)-op_{2,1}(n,k+1).$
For $k\geq 1$, the generating function of $op_{2,1}(n,k+1)$ is
$\begin{split}\sum_{n=1}^{\infty}op_{2,1}(n,k+1)&=(-q;q)_{\infty}\sum_{j=0}^{\infty}\frac{q^{1+3+\cdots+(2k+4j+1)}(1-q^{2k+4j+3})}{(q;q)_{\infty}}\\ &=\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).\end{split}$ (2.6)
Then, the following truncated version of (1.2) immediately follows from (2.5) and (2.6).
Corollary 2.6.
For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).$
Combining Theorem 1.3, Theorem 1.6 and Corollary 2.5, we can get the following corollary.
Corollary 2.7.
For $n,k\geq 1$,
$op_{2,1}(n,k+1)=\overline{M}_{k}(n)/{2},$ (2.7)
and
$op_{2,1}(n,k)-op_{2,1}(n,k+1)=\overline{N}_{k}(n).$
Then, we can get (1.7) by using (1.1) with $q\rightarrow q^{\ell}$, (2.6), (2.7), and Euler’s partition identity [1, (1.2.5)]:
$\frac{1}{(q;q^{2})_{\infty}}=(-q;q)_{\infty}.$
Note that $\overline{M}_{0}(n)=\overline{p}(n)$ for $n\geq 1$, then by Theorem 2.3, we have
$op_{2,1}(n,1)=\overline{M}_{0}(n)/{2}.$
Combining with Corollary 2.7, we get the following result.
Corollary 2.8.
For $k\geq 1$,
$\sum_{n=1}^{\infty}\left(\overline{M}_{k-1}(n)-\overline{M}_{k}(n)\right)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
This article is organized as follows. We first give a combinatorial proof of (1.6) in Section 3. Then, we show Theorems 2.3 and 2.4 in Section 4. Finally, we give an analytic proof of Corollary 2.8 in Section 5.
3 Combinatorial proof of (1.6)
Clearly, in order to prove (1.6), it is equivalent to showing that for $n\geq 1$,
$\overline{p}(n)/2\leq\overline{p}(n-1),\text{ with strict inequality if }n\geq 4.$ (3.1)
To do this, we introduce three sets of overpartitions. Bear in mind that the parts in an overpartition are ordered as in (1.4).
For $n\geq 1$, let $\mathcal{A}(n)$ be the set of overpartitions of $n$ such that the smallest part is non-overlined. For example, we have
$\mathcal{A}(4)=\{(4),(3,1),(\overline{3},1),(2,2),(2,1,1),(\overline{2},1,1),(1,1,1,1)\}.$
For an overpartition $\pi$ in $\mathcal{A}(n)$, if we change the smallest part to an overlined part, then we get an overpartition of $n$ such that the smallest part is overlined, and vice versa. This implies that the number of overpartitions in $\mathcal{A}(n)$ is $\overline{p}(n)/2$.
For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part lager than $1$ in $\lambda$ is greater than or equal to two plus the number of parts $1$ in $\lambda$. For example, we have
$\mathcal{B}(3)=\{(3),(\overline{3}),(2,1),(\overline{2},1),(2,\overline{1}),(1,1,1),(1,1,\overline{1})\}.$
For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part larger than $1$ in $\lambda$ is less than two plus the number of parts $1$ in $\lambda$. For example, we have
$\mathcal{C}(3)=\{(\overline{2},\overline{1})\}.$
To give a combinatorial proof of (3.1), it suffices to show the following theorem.
Theorem 3.1.
There exists a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$ and $\mathcal{C}(n-1)$ is nonempty for $n\geq 4$.
Proof. For $n\geq 1$, let $\pi=(\pi_{1},\ldots,\pi_{m-1},\pi_{m})$ be an overpartition in $\mathcal{A}(n)$. By definition, we know that $\pi_{m}$ is a non-overlined part, which implies that $\overline{1}$ does not appear in $\pi$. Assume that $\pi_{m}=t$, we consider the following two cases.
Case 1: $t=1$. In this case, we set $\lambda=(\pi_{1},\ldots,\pi_{m-1})$.
Case 2: $t\geq 2$. In this case, we set
$\lambda=(\pi_{1},\ldots,\pi_{m-1},\underbrace{1,\ldots,1}_{(t-2)^{\prime}\text{s}},\overline{1}).$
In either case, we get an overpartition $\lambda$ in $\mathcal{B}(n-1)$. Obviously, the process above is reversible. Now, we have built a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$.
It remains to show that $\mathcal{C}(n-1)$ is nonempty for $n\geq 4$. It can be checked that for $n\geq 4$,
$(\overline{2},\underbrace{1,\ldots,1}_{(n-4)^{\prime}\text{s}},\overline{1})$
is an overpartition in $\mathcal{C}(n-1)$. This completes the proof.
4 Proofs of Theorems 2.3 and 2.4
The objective of this section is to give the proofs of Theorems 2.3 and 2.4. To do this, we need the following lemma.
Lemma 4.1.
For $n,j\geq 1$,
$\overline{p}(n-j^{2})=op_{2,1}(n,j)+op_{2,1}(n,j+1).$
Proof. For an overpartition $\pi$ of $n-j^{2}$, if we add the parts $1,3,\ldots,2j-1$ into $\pi$, then we get an overpartition $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$, and vice versa. It implies that $\overline{p}(n-j^{2})$ equals the number of overpartitions $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$.
On the other hand, it follows from definition that $op_{2,1}(n,j)+op_{2,1}(n,j+1)$ is also the number of overpartitions $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$. This completes the proof.
Then, we give a proof of Theorem 2.3 with the aid of Lemma 4.1.
Proof of Theorem 2.3. Divide both sides of (1.2) by $\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}$, we can get
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}\right)=1.$ (4.1)
Comparing the coefficients of $q^{n}$ on the both sides of (4.1), we have
$\overline{p}(n)+2\sum_{j=1}^{\infty}(-1)^{j}\overline{p}(n-j^{2})=0.$
Combining with Lemma 4.1, we get
$\overline{p}(n)=-2\sum_{j=1}^{\infty}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)=2op_{2,1}(n,1).$
The proof is complete.
We conclude this section with a proof of Theorem 2.4.
Proof of Theorem 2.4. We first show (2.3). There are the following two cases.
Case 1: $k\geq m\geq 1$. In this case, we have $\text{min}\{|m|,k\}=|m|=m$, $\text{min}\{|m|,|k|\}=|m|=m$ and $\text{max}\{|m|,|k|\}=|k|=k$. It follows from Lemma 4.1 that
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})$
$=(-1)^{m}\sum_{j=m}^{k}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)$
$=(-1)^{m}\left((-1)^{m}op_{2,1}(n,m)+(-1)^{k}op_{2,1}(n,k+1)\right)$
$=op_{2,1}(n,m)+(-1)^{m+k}op_{2,1}(n,k+1),$
which agrees with (2.3) for $k\geq m\geq 1$.
Case 2: $m\leq k\leq-1$. In this case, we have $\text{min}\{|m|,k\}=k$, $\text{min}\{|m|,|k|\}=|k|=-k$ and $\text{max}\{|m|,|k|\}=|m|=-m$. Again by Lemma 4.1, we have
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})$
$=(-1)^{k}\sum_{j=-k}^{-m}(-1)^{j}\overline{p}(n-j^{2})$
$=(-1)^{k}\sum_{j=-k}^{-m}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)$
$=(-1)^{k}\left((-1)^{-k}op_{2,1}(n,-k)+(-1)^{-m}op_{2,1}(n,-m+1)\right)$
$=op_{2,1}(n,-k)+(-1)^{m+k}op_{2,1}(n,-m+1).$
We arrive at (2.3) for $m\leq k\leq-1$.
Now, we have completed the proof of (2.3). Then, we proceed to show (2.4). Under the condition that $m\leq k$ and $mk\leq 0$, we have $m\leq 0\leq k$. Clearly, we have
$\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=\sum_{j=m}^{-1}(-1)^{j}\overline{p}(n-j^{2})+\overline{p}(n)+\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2}).$ (4.2)
In view of (2.3), we get
$\sum_{j=m}^{-1}(-1)^{j}\overline{p}(n-j^{2})=-\left(op_{2,1}(n,1)+(-1)^{m+1}op_{2,1}(n,-m+1)\right),$ (4.3)
and
$\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})=-\left(op_{2,1}(n,1)+(-1)^{k+1}op_{2,1}(n,k+1)\right).$ (4.4)
Substituting (4.3) and (4.4) into (4.2) and combining with Theorem 2.3, we have
$\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=(-1)^{m}op_{2,1}(n,-m+1)+(-1)^{k}op_{2,1}(n,k+1).$ (4.5)
Then, we consider the following two cases.
Case 1: $k\geq-m\geq 0$. In this case, we have $\text{min}\{|m|,k\}=|m|=-m$, $\text{min}\{|m|,|k|\}=|m|=-m$ and $\text{max}\{|m|,|k|\}=|k|=k$. Appealing to (4.5), we get
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,-m+1)+(-1)^{m+k}op_{2,1}(n,k+1).$
So, (2.4) is valid for $k\geq-m\geq 0$.
Case 2: $m>k\geq 0$. In this case, we have $\text{min}\{|m|,k\}=k$, $\text{min}\{|m|,|k|\}=|k|=k$ and $\text{max}\{|m|,|k|\}=|m|=-m$. Using (4.5), we get
$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,k+1)+(-1)^{m+k}op_{2,1}(n,-m+1),$
and thus (2.4) is satisfied for $m>k\geq 0$. This completes the proof.
5 Analytic proof of Corollary 2.8
In this section, we aim to give an analytic proof of Corollary 2.8. Note that for $k\geq 0$, the generating function of $\overline{M}_{k}(n)$ is
$\sum_{n=1}^{\infty}\overline{M}_{k}(n)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$
so it remains to show
$\begin{split}&\quad\frac{(-q;q)_{k-1}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\\ &=\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.\end{split}$ (5.1)
Dividing both sides of (5.1) by $\frac{(-q;q)_{k-1}}{(q;q)_{k}}$, we find that it is equivalent to
$\begin{split}&\quad(1-q^{k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\\ &=(1-q^{2k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.\end{split}$ (5.2)
Then, we proceed to show (5.2).
Proof of (5.2). Note that
$\quad(1+q^{k})q^{(k+1)(k+j+1)}$
$=q^{(k+1)(k+j+1)}+q^{(k+1)(k+j+1)+k}$
$=q^{k(k+j+1)}(q^{k+j+1}-1+1)+q^{k(k+j+2)}(q^{k+j+1}+1-1)$
$=q^{k(k+j+1)}(1-q^{k})-q^{k(k+j+1)}(1-q^{k+j+1})+q^{k(k+j+2)}(1+q^{k+j+1}),$
so we get
$\quad(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=\sum_{j=0}^{\infty}\left((1-q^{k})\frac{q^{k(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\frac{q^{k(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+2};q)_{\infty}}\right$.
$\qquad\qquad\qquad\left.+\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\right)$
$=(1-q^{k})\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$\qquad\qquad\qquad+\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
Then, we have
$\quad(1-q^{k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=(1-q^{k})\frac{q^{k^{2}}(-q^{k+1};q)_{\infty}}{(q^{k};q)_{\infty}}+\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=(1-q^{2k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
We arrive at (5.2). This completes the proof.